OpenStudy (anonymous):

Lef f(x)=-2x+7 and g(x)=-6x+3. find F(closed o) g and state it's domain

OpenStudy (anonymous):

i mean let =/

OpenStudy (anonymous):

F(closed o) g= f(g(x))= f(-6x+3)= -2(-6x+3)+7= *12x+1*

OpenStudy (anonymous):

possible answers are ...8x-4 ..-4x-10 ...8x-10 ...-4x-4